Determine a, h and k, of the translated function `g(x)`based from the description, then write the equation of the quadratic function in vertex form.
1.) The graph of f(x)=2x^2 is reflected across the x-axis, shifted 6 units to the right and 1 unit upward.
Identify a, h and k: (give the values only no need to explain)
Equation for #1:
2.) The graph of f(x)=x^2 is shifted half unit down and half unit to the right.
Identify a, h and k:
equation for #2:
3.) The parent function ff(x)=2x^2 is reflected across the x-axis and translated 9 units right and 5 units up.
Identify a, h and k:
equation for #3:

Answer by CPhill(1959)

Here are the solutions:
**1.**
* a = -2
* h = 6
* k = 1
Equation: g(x) = -2(x - 6)² + 1
**2.**
* a = 1
* h = 1/2
* k = -1/2
Equation: g(x) = (x - 1/2)² - 1/2
**3.**
* a = -2
* h = 9
* k = 5
Equation: g(x) = -2(x - 9)² + 5